The area of Beth's rectangular garden is(x2+13x+42)ft.2.the with is(x+6)ft.
what is the length of the garden?
Find the perimeter in term x.

Answer by dabanfield(803)

The area of Beth's rectangular garden is(x2+13x+42)ft.2.the with is(x+6)ft.
what is the length of the garden?
Area = length*width so
length = area/width
So in this case:
length = (x^2+13x+42)/(x+6) or
length = [(x+6)*(x+7)]/(x+6)
length = x+7
The perimeter P of a rectangle is 2*length + 2*width
In this case:
P = 2*(x+7) + 2*(x+6)